OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. \(-9-3+3+9+...+81\)

OpenStudy (michele_laino):

we have an arithmetic sequence, whose firs term is -9, and the constant is: d= -3-(-9)=3-(-3)=...

OpenStudy (anonymous):

Would it be \[\huge \sum_{n=0}^{\infty}=(-9+6n)\] because the common difference is 6?

OpenStudy (anonymous):

Wait, it should 15 not infinity. :D

geerky42 (geerky42):

That's correct.